Circular No. 6052

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERSEID METEORS 1994 P. Brown, University of Western Ontario, writes: “In spite of poor

weather across North America, several observers have reported significant Perseid activity. Reports from European and North American observers in the time interval Aug. 11.88–12.40 UT suggest that activity was not then significantly enhanced. J. Rendtel, I. Rendtel, A. Knoefel, and D. Holman (International Meteor Organization), observing from the Nevada-California border, northwest of Reno under excellent conditions, recorded steadily increasing rates from Aug. 12.39 to 12.46 with a maximum zenithal hourly rate (ZHR) in the range 250–300 at Aug. 12.45. These observers noted a steep decline in activity beginning around Aug. 12.46 and ended their observations in twilight near Aug. 12.48 (when ZHR ∼ 75). The observers noted several fireballs in the interval Aug. 12.44–12.46 and commented on the rapid fluctuation in meteor rates during this interval.”

X-RAY NOVA IN SCORPIUS M. Della Valle, Department of Astronomy, Universitá di Padova, re-

ports: “Preliminary analysis of flux-calibrated spectrograms (range 360-900 nm, resolution ∼ 0.8 nm) obtained on Aug. 12.1 UT at the European Southern Observatory’s 1.5-m telescope (+ Boller & Chivens spectrograph) confirms the candidate reported by Bailyn et al. (IAUC 6050) to be the optical counterpart of GRO J1655–40 (IAUC 6046). The spectrum exhibits prominent, broad Balmer lines (Hα, FWHM ∼ 4.7 nm; Hβ, FWHM ∼ 3 nm, corresponding to velocities of ∼ 2000 km/s) superimposed on a relatively red continuum. Supplementary emission lines are measured at 668.2 (He II mult. 7), 468.4 (He II mult. 1), and 463.6 nm (N III).”

D. Campbell-Wilson and R. Hunstead, University of Sydney, report the results of radio observations of the field of GRO J1655-40, using the Molonglo Observatory Synthesis Telescope at 843 MHz: “A strong variable radio source has been detected within 5′′ of the position of the optical nova reported by Bailyn et al. (IAUC 6050). A 3-hr synthesis observation obtained on Aug. 6.60 UT (mid-exposure) gave a flux density of 370 mJy (± 6 percent), and a second 3-hr observation on Aug. 11.59 gave 770 mJy (± 6 percent). More recent observations confirm that the flux increase is continuing, with a level of around 900 mJy being recorded at Aug. 12.33. Monitoring at 843 MHz is continuing, and observations at other frequencies are urgently needed.”

1994 August 12 Daniel W. E. Green